### v1

@@ 4.3 Performance analysis
$N$ denotes the number of loop iterations, $P$ is the number of OpenMP threads, and $\phi=1.618$ as defined by Mohammed et al. (2022).
(a) $(n1,n2,n3)=(100,400,400)$
(b) $(n1,n2,n3)=(200,400,400)$
(c) $(n1,n2,n3)=(400,400,400)$
Figure 4: Single-shot FWI runtime using the proposed autotuning compared to the default OpenMP static and guided schedulers, and the OpenMP dynamic using the chunk size proposed by Mohammed et al. (2022). This set of experiments was performed on NPAD machine for three input sizes, $(n1,n2,n3)=$ (a) $(100,400,400)$, (b) $(200,400,400)$, and (c) $(400,400,400)$. Each point is a median of at least five runs.
As Fig. 4 shows, our proposed method presented the lowest runtimes in all scenarios tested in this set of experiments. In comparison to the case where the chunk size is the one defined by Mohammed et al. (2022), our autotuning performed about $4\%$ faster for the three input sizes we tested.
@@ 5 Related Work
Autotuning has been widely studied. Some authors present methods to adjust workflows automatically. HoseinyFarahabady et al. (2020) proposed an autotuning method to improve the data flow of stream processing applications based on continuously collected runtime information (e.g., available resources and incoming traffic rate). Shu et al. (2021) proposed the Component-based Ensemble Active Learning (CEAL). This autotuning method combines machine learning techniques with knowledge of the in-situ workflow structure to enable automated workflow configuration with limited performance measurements. Our proposal aims to autotune applications, unlike HoseinyFarahabady et al. (2020); Shu et al. (2021).
@@
Application autotune has been exploited from several perspectives. Padoin et al. (2014, 2017) employ load balancing techniques to control processor frequency, improving the energy efficiency of parallel applications on multicore systems. Both works use processor frequency tuning techniques to slow down less loaded cores, saving energy. Rocha et al. (2020) presents AtTune, a heuristic-based framework to optimize the parallel execution of applications, tuning the number of processes/threads and CPU frequency levels. Bez (2021) proposes to use machine learning techniques to automatically detect I/O access patterns at runtime for applications in High-Performance Computing (HPC) platforms. Based on those patterns, it is possible to dynamically adjust I/O parameters and allocate resources more efficiently to improve the system’s overall performance. Bağbaba et al. (2021) presented an autotuning approach based on I/O monitoring and predictive modeling. This approach can find a set of I/O parameter values in a given system and application use case. This autotuning is applied to parallel I/O stacks in distributed systems using the MPI-IO ROMIO library. Robert et al. (2020) developed an optimization tool called SHAMAN, which provides a ready-to-use web application to perform black-box autotuning of custom computer components in a distributed system for a specific application executed through the Slurm workload manager. Different from Padoin et al. (2014, 2017); Rocha et al. (2020); Bez (2021); Bağbaba et al. (2021); Robert et al. (2020), our autotuning works on code-level optimization.
@@
Several works proposed approaches for code-level optimization autotuning. Kale et al. (2014) proposed to improve scheduling by combining static and dynamic approaches in symmetric multiprocessing (SMP) machines, emphasizing spatial locality. Different from Kale et al. (2014), we propose an autotuning for any parallel application using OpenMP.
@@
Katagiri et al. (2014, 2015) presented ppOpen-AT, a framework for directive-driven code optimization such as loop split and loop fusion. Menon et al. (2020) presented HiPerBOt, a configuration selection framework based on Bayesian optimization, to identify application and platform-level parameters that result in high-performance configurations. Liu et al. (2021) present GPTune, an autotuning based on multitask learning and Bayesian optimization suitable for tuning exascale application codes. Kimovski et al. (2021) developed a machine learning-based framework to optimize exascale applications. The AutoTuner can adjust the application’s parameters to improve performance when detecting an anomaly. Roy et al. (2021) presented Bliss, an algorithm that uses Bayesian optimization to find the near-optimal parameter configuration. Bliss must run the application a few times to determine the cost function. Milani (2020) proposed policies to choose the best among user-defined code versions automatically. In common, Katagiri et al. (2014, 2015); Menon et al. (2020); Liu et al. (2021); Kimovski et al. (2021); Roy et al. (2021); Milani (2020) need to run the application under a variety of different configurations to learn the correlation between an application’s parameters and its performance, which can be prohibitive for computationally expensive applications such as FWI. We propose using the time to compute the first iteration of the laplacian as an estimative of the time of the FWI.
@@
Sakurai et al. (2020) proposed two autotuning functions for heterogeneous computing architectures. The first function may perform loop transformations to adjust the position of directives in OpenMP loops automatically. The second function dynamically changes the number of threads. Bak et al. (2018) introduced a load-balancing method based on integrating Charm++ and OpenMP to distribute user-created tasks dynamically. Kruse et al. (2020) addressed compiler performance optimization through a loop transformation autotuning. Their proposal is based on search trees to explore the loop transformation space and uses pragmas inserted in the source code to apply these transformations. Our proposal aims to optimize parallel loops scheduling, unlike Sakurai et al. (2020); Bak et al. (2018); Kruse et al. (2020).
@@
Many authors proposed autotuning methods suitable for optimizing parallel loops scheduling. Dutta et al. (2022) presented preliminary work on using deep machine learning to identify loop patterns and adjust loop parameters with OpenMP. As Dutta et al. (2022) use machine learning, it is necessary to train it before execution. Wood et al. (2021) introduced Artemis, an online framework that dynamically tunes the execution of parallel regions by training optimizing models. Our proposed method requires no training, unlike Dutta et al. (2022); Wood et al. (2021).
@@
Seiferth et al. (2020) introduced an approach that autotune applications by classifying optimization approaches based on a performance analytical model. Andreolli et al. (2014, 2015) brought an approach to automatically tune seismic applications by compiling and executing each set of parameters chosen by a genetic algorithm, including load block size and compilation flags. Seiferth et al. (2020); Andreolli et al. (2014, 2015) methods perform autotuning statically. That way, they may encounter a system state different from the runtime, as aspects such as memory availability, resource usage and access, workload, and cores’ load may vary over time. On the other hand, our method performs autotuning dynamically at runtime, allowing it to adjust the application for the system’s current state.
@@
Dennis Booth and Allen Lane (2022) presented iCh, an autotuning method for OpenMP loop schedulers (guided and dynamic) that uses low-cost heuristics to reduce load imbalance. iCh uses on-demand work-stealing as its primary load-balancing mechanism and a distribution system with local queues for threads. By using work-stealing, iCh may have extra parallel overhead as it has to manage queues of tasks and thread synchronization. On the contrary, our proposal uses the OpenMP dynamic schedule.
@@
Assis et al. (2020) proposed using PATSMA with CSA to optimize the chunk size of parallel loops of a reverse time migration (RTM). Here, we expand Assis et al. (2020)’s approach to make it suitable for an FWI. FWI poses an extra challenge to autotuning methods as it has higher algorithmic complexity and a potentially different input for each iteration. This paper also increases reproducibility as tests were also performed in cloud instances. Finally, we present comparative tests with a state-of-the-art autotuning method.
@@ 6 Conclusion
This work proposed an autotuning strategy based on the Parameter Auto Tuning for Shared Memory Algorithms (PATSMA) to find an ideal chunk size for a dynamic OpenMP scheduler. This approach was designed to work in different computing environments, regardless of parameters such as number of threads, processors, and amount of memory. We applied the proposed mechanism to a Full Waveform Inversion (FWI) algorithm, reducing its execution time.
Experiments varying the input size and in different computing environments showed that the proposed autotuning outperforms the default OpenMP static and guided schedulers in all tested scenarios. For those experiments, the proposed strategy achieved speedups of up to $70.46\%$. Those results demonstrate that our method adapts well to different problem sizes and computing environments, outperforming the OpenMP standard schedulers.
Another set of experiments applied to an FWI showed the performance of the proposed method compared to the OpenMP static and guided schedulers for different numbers of seismic shots. In this set of tests, the proposed autotuning also outperformed the OpenMP schedulers in tested scenarios, reaching up to $7.98\%$ speedup. In this set of tests, it was observed that the speedup remained stable because the overhead did not increase with the increase in the number of shots. It is possible to conclude that the chunk size adjusted for the first shot can be reused for the following shots and iterations of FWI. In all cases, the overhead was below $1.2\%$. The results also show that the proposed method is scalable, presenting higher speedups for tests with larger input sizes and on architectures with more cores.
Finally, we have compared the execution times of an FWI using the OpenMP dynamic with the chunk size determined by our autotuning and by the formula proposed by Mohammed et al. (2022). For different input sizes, our method performed about $4\%$

### v2

@@ 4.3 Performance analysis
$N$ denotes the number of loop iterations, $P$ is the number of OpenMP threads, and $\phi=1.618$. We ran this set of experiments in SD and NPAD machines.
Fig. 5 shows the FWI runtimes measured in the SD machine. For all input sizes, the median runtime of the FWI using the OpenMP dynamic scheduler with the chunk size defined by Mohammed et al. (2022) was lower than using the OpenMP default static and guided schedulers. Also, the FWI using our proposed method presented the lowest runtimes in all scenarios tested in this set of experiments being $3.75\%$, $3.88\%$, and $3.86\%$ faster than the FWI using the OpenMP dynamic scheduler with the chunk size defined by Mohammed et al. (2022) for $n1=100$, $n1=200$, and $n1=400$, respectively.
Figure 5: Single-shot FWI runtime using the proposed auto-tuning compared to the default OpenMP static and guided schedulers, and the OpenMP dynamic, static, and guided using the chunk size proposed by Mohammed et al. (2022) (marked with *). This set of experiments was performed on the SD machine for three input sizes, $(n1,n2,n3)=$ (a) $(100,400,400)$, (b) $(200,400,400)$, and (c) $(400,400,400)$. Each point is a median of five runs.
Fig. 6 shows the FWI runtimes measured in the NPAD machine. For all input sizes, the median runtime of the FWI with the chunk size defined by Mohammed et al. (2022) was higher than using the OpenMP default static and guided schedulers. Also, the FWI using our proposed method presented the lowest runtimes in all scenarios tested in this set of experiments being $2.07\%$, $3.05\%$, and $5.12\%$ faster than the fastest FWI using the default OpenMP dynamic or static schedulers for $n1=100$, $n1=200$, and $n1=400$, respectively.
Figure 6: Single-shot FWI runtime using the proposed auto-tuning compared to the default OpenMP static and guided schedulers, and the OpenMP dynamic, static, and guided using the chunk size proposed by Mohammed et al. (2022) (marked with *). This set of experiments was performed on the SD machine for three input sizes, $(n1,n2,n3)=$ (a) $(100,400,400)$, (b) $(200,400,400)$, and (c) $(400,400,400)$. Each point is a median of five runs.
@@ 5 Related Work
Auto-tuning has been widely studied. Some authors present methods to adjust workflows automatically. HoseinyFarahabady et al. (2020) proposed an auto-tuning method to improve the data flow of stream processing applications based on continuously collected runtime information (e.g., available resources and incoming traffic rate). Shu et al. (2021) proposed the Component-based Ensemble Active Learning (CEAL). This auto-tuning method combines machine learning techniques with knowledge of the in-situ workflow structure to enable automated workflow configuration with limited performance measurements. Our proposal aims to autotune applications, unlike HoseinyFarahabady et al. (2020); Shu et al. (2021).
@@
Application autotune has been exploited from several perspectives. Padoin et al. (2014, 2017) employ load balancing techniques to control processor frequency, improving the energy efficiency of parallel applications on multicore systems. Both works use processor frequency tuning techniques to slow down less loaded cores, saving energy. Rocha et al. (2020) presents AtTune, a heuristic-based framework to optimize the parallel execution of applications, tuning the number of processes/threads and CPU frequency levels. Bez (2021) proposes to use machine learning techniques to automatically detect I/O access patterns at runtime for applications in High-Performance Computing (HPC) platforms. Based on those patterns, it is possible to dynamically adjust I/O parameters and allocate resources more efficiently to improve the system’s overall performance. Bağbaba et al. (2021) presented an auto-tuning approach based on I/O monitoring and predictive modeling. This approach can find a set of I/O parameter values in a given system and application use case. This auto-tuning is applied to parallel I/O stacks in distributed systems using the MPI-IO ROMIO library. Robert et al. (2020) developed an optimization tool called SHAMAN, which provides a ready-to-use web application to perform black-box auto-tuning of custom computer components in a distributed system for a specific application executed through the Slurm workload manager. Different from Padoin et al. (2014, 2017); Rocha et al. (2020); Bez (2021); Bağbaba et al. (2021); Robert et al. (2020), our auto-tuning works on code-level optimization.
@@
Several works proposed approaches for code-level optimization auto-tuning. Kale et al. (2014) proposed to improve scheduling by combining static and dynamic approaches in symmetric multiprocessing (SMP) machines, emphasizing spatial locality. Different from Kale et al. (2014), we propose an auto-tuning for any parallel application using OpenMP.
@@
Katagiri et al. (2014, 2015) presented ppOpen-AT, a framework for directive-driven code optimization such as loop split and loop fusion. Menon et al. (2020) presented HiPerBOt, a configuration selection framework based on Bayesian optimization, to identify application and platform-level parameters that result in high-performance configurations. Liu et al. (2021) present GPTune, an auto-tuning based on multitask learning and Bayesian optimization suitable for tuning exascale application codes. Kimovski et al. (2021) developed a machine learning-based framework to optimize exascale applications. The AutoTuner can adjust the application’s parameters to improve performance when detecting an anomaly. Roy et al. (2021) presented Bliss, an algorithm that uses Bayesian optimization to find the near-optimal parameter configuration. Bliss must run the application a few times to determine the cost function. Milani (2020) proposed policies to choose the best among user-defined code versions automatically. In common, Katagiri et al. (2014, 2015); Menon et al. (2020); Liu et al. (2021); Kimovski et al. (2021); Roy et al. (2021); Milani (2020) need to run the application under a variety of different configurations to learn the correlation between an application’s parameters and its performance, which can be prohibitive for computationally expensive applications such as FWI. We propose using the time to compute the first iteration of the laplacian as an estimative of the time of the FWI.
@@
Sakurai et al. (2020) proposed two auto-tuning functions for heterogeneous computing architectures. The first function may perform loop transformations to adjust the position of directives in OpenMP loops automatically. The second function dynamically changes the number of threads. Bak et al. (2018) introduced a load-balancing method based on integrating Charm++ and OpenMP to distribute user-created tasks dynamically. Kruse et al. (2020) addressed compiler performance optimization through a loop transformation auto-tuning. Their proposal is based on search trees to explore the loop transformation space and uses pragmas inserted in the source code to apply these transformations. Our proposal aims to optimize parallel loops scheduling, unlike Sakurai et al. (2020); Bak et al. (2018); Kruse et al. (2020).
@@
Many authors proposed auto-tuning methods suitable for optimizing parallel loops scheduling. Dutta et al. (2022) presented preliminary work on using deep machine learning to identify loop patterns and adjust loop parameters with OpenMP. As Dutta et al. (2022) use machine learning, it is necessary to train it before execution. Wood et al. (2021) introduced Artemis, an online framework that dynamically tunes the execution of parallel regions by training optimizing models. Our proposed method requires no training, unlike Dutta et al. (2022); Wood et al. (2021).
@@
Seiferth et al. (2020) introduced an approach that autotunes applications by classifying optimization approaches based on a performance analytical model. Andreolli et al. (2014, 2015) brought an approach to automatically tune seismic applications by compiling and executing each set of parameters chosen by a genetic algorithm, including load chunk size and compilation flags. Seiferth et al. (2020); Andreolli et al. (2014, 2015) methods perform auto-tuning statically. That way, they may encounter a system state different from the runtime, as aspects such as memory availability, resource usage and access, workload, and cores’ load may vary over time. On the other hand, our method performs auto-tuning dynamically at runtime, allowing it to adjust the application for the system’s current state.
@@
Dennis Booth and Allen Lane (2022) presented iCh, an auto-tuning method for OpenMP loop schedulers (guided and dynamic) that uses low-cost heuristics to reduce load imbalance. iCh uses on-demand work-stealing as its primary load-balancing mechanism and a distribution system with local queues for threads. By using work-stealing, iCh may have extra parallel overhead as it has to manage queues of tasks and thread synchronization. On the contrary, our proposal uses the OpenMP dynamic schedule.
@@
Assis et al. (2020) proposed using PATSMA with CSA to optimize the chunk size of parallel loops of a reverse time migration (RTM). Here, we expand Assis et al. (2020)’s approach to make it suitable for an FWI. FWI poses an extra challenge to auto-tuning methods as it has higher algorithmic complexity and a potentially different input for each iteration. This paper also increases reproducibility as tests were also performed in cloud instances. Finally, we present comparative tests with a state-of-the-art auto-tuning method.
@@ 6 Conclusion
This work proposed an auto-tuning strategy based on the Parameter Auto-Tuning for Shared Memory Algorithms (PATSMA) to find an ideal chunk size for a dynamic OpenMP scheduler. This approach was designed to work in different computing environments, regardless of parameters such as number of threads, processors, and amount of memory. We applied the proposed mechanism to a Full Waveform Inversion (FWI) algorithm, reducing its execution time.
Experiments varying the input size and in different computing environments showed that the proposed auto-tuning outperforms the default OpenMP static and guided schedulers in all tested scenarios. For those experiments, the proposed strategy achieved speedups of up to $70.46\%$. Those results demonstrate that our method adapts well to different problem sizes and computing environments, outperforming the OpenMP standard schedulers.
Another set of experiments applied to an FWI showed the performance of the proposed method compared to the OpenMP static and guided schedulers for different numbers of seismic shots. In this set of tests, the proposed auto-tuning also outperformed the OpenMP schedulers in tested scenarios, reaching up to $14.96\%$ speedup. In this set of tests, it was observed that the speedup remained stable because the overhead did not increase with the increase in the number of shots. It is possible to conclude that the chunk size adjusted for the first shot can be reused for the following shots and iterations of FWI. In all cases, the overhead was below $1.2\%$. The results also show that the proposed method is scalable, presenting higher speedups for tests with larger input sizes and on architectures with more cores.
Finally, we have compared the execution times of an FWI using our proposed auto-tuning, the default OpenMP static and guided schedulers, and the OpenMP static, dynamic and guided schedulers with the chunk size determined by the formula proposed by Mohammed et al. (2022). Those tests were performed in two machines for three input sizes. Our method outperformed the schedulers tested in all scenarios, reaching speedups from $2.07\%$ to $34.95\%$.
Computer code availability
Code essential to the analysis in this paper is available on GitLab at the following public repositories: (1) https://gitlab.com/lappsufrn/seismic/ufrn-fwi/mamute Full Waveform Inversion code in a repository called Mamute and (2) https://gitlab.com/lappsufrn/patsma PATSMA code. All codes utilized in this paper were written in C++ language and are open source under the MIT License.